Two Current Carrying Wire Along To X Axis In Arbitary Direction.current is I & -I respectively.So Can You Give Me The All Possible Condition , SAGAR SINGH

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

\begin{figure} \begin{center} \leavevmode \epsfxsize=2.8 in \epsfbox{/export/home/fyde/randy/figs/fig19-9.eps}\end{center}\end{figure}

Wire #2 will experience a magnetic field  due to wire #1 given by

 


B1 = $\displaystyle{\frac{\mu_0I_1}{2\pi r}}$

in the direction indicated, and hence will experience a force per unit length given by 


$\displaystyle{\frac{F_2}{l}}$ = $\displaystyle{\frac{\mu_0I_1I_2}{2\pi d}}$.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com